You want to be able to withdraw $20,000 each year for 15 years. Your account earns 7% interest.

a) How much do you need in your account at the beginning? $

b) How much total money will you pull out of the account? $

c) How much of that money is interest? $
Answer #1

Amount needed at the beginning = Annual Withdrawal*Present value annuity factor

= 20,000*PVAF(7%, 15 years)

= 20,000*9.1079

= $182,158

b)Total money withdrawn = 20,000*15 = $300,000

c)Interest amount = 300,000-182,158 = $117,842